# Bound States and Particle Production by Breather-Type Background Field Configurations

### v1

@@ 6 Conclusions
A key finding is that although fermion wave functions initially appear bound to the domain walls, over time they do not remain localized, and indeed they do not remain significantly more localized than arbitrary wave functions that are not closely approximated by bound states of the static forms of the potentials. In either case, the oscillating backgrounds leads to the propagation of fermionic states outward at speeds up to the speed of light, as demonstrated by the cone-shaped shock fronts seen in the graphs (i.e. figure 1 for $g=0.10$, figure 2 for $g=0.25$, and figure 3 for $g=0.50$). These graphs clearly depict the expansion of the fermion density beyond the domain wall region, showing the spread of fermion wave functions across space.
Numerical simulations show that even though some particle density remains localized near the initial positions of the domain walls after one oscillation period (visible as partial recurrences in figure 5), the overall trend is a wave of flux moving toward spatial infinity. This feature underscores the limitations of approximating such systems as fully integrable, as the outward particle flux becomes a significant factor in the system’s long-term behavior.
Further computational challenges arise due to the difficulty in imposing stable boundary conditions, particularly in light of the fermionic modes propagating outward. Attempts to impose radiation boundary conditions encountered issues, as damping methods applied to the Dirac equation led to standing waves rebounding between the integration domain boundaries. This behavior is illustrated by the numerical instabilities at the boundaries that arise over extended integration times, as discussed in the paper.
@@
Overall, the study concludes that the sine-Gordon-like breather background, when coupled to fermions, does not support steady-state bound fermion solutions. Instead, the oscillations induce fermion particle production and flux propagation, suggesting that the fermion-soliton system does not maintain the integrability and stability expected from the purely bosonic sine-Gordon theory.

### v2

@@ 6 Conclusions
A key finding is that although fermion wave functions initially appear bound to the domain walls, over time they do not remain localized, and indeed they do not remain significantly more localized than arbitrary wave functions that are not closely approximated by bound states of the static forms of the potentials. In either case, the oscillating backgrounds leads to the propagation of fermionic states outward at speeds up to the speed of light, as demonstrated by the cone-shaped shock fronts seen in the graphs (i.e. figure 1 for $g=0.10$, figure 2 for $g=0.25$, and figure 3 for $g=0.50$). These graphs clearly depict the expansion of the fermionic excitations beyond the domain wall regions, showing the spread of interacting fermion wave functions across space. This is rather unlike the behavior often seen in the purely bosonic sectors of higher-order non-integrable models [44, 45], in which the effective potentials between the solitary waves in $1+1$ dimensions [12, 46] tend to remain well localized, even when the individual kinks and antikinks possess fat tails [13].
Numerical simulations show that even though some particle density remains localized near the initial positions of the domain walls after one oscillation period (visible as partial recurrences in figure 5), the overall trend is a wave of flux moving toward spatial infinity. This feature underscores the limitations of approximating such systems as fully integrable, as the outward particle flux becomes a significant factor in the system’s long-term behavior. Moreover, after a sharp falloff in the fermion field amplitude at small times, the oscillating fermion fields do not damp to zero at large times. Instead, their amplitudes remain finite and comparable throughout the light cone interiors, as the motion of the kink-antikink pair continues to pump energy into the fermion sector.
Further computational challenges arise due to the difficulty in imposing stable boundary conditions, particularly in light of the fermionic modes propagating outward. Attempts to impose radiation-like boundary conditions encountered confounding issues, as damping methods applied to the Dirac equation led to standing waves rebounding between the integration domain boundaries. This behavior is illustrated by the numerical instabilities at the boundaries that arise over extended integration times, as discussed in section 5.
@@
Overall, this analysis concludes that the sine-Gordon-like breather background, when coupled to fermions, does not generally support steady-state bound fermion solutions. Instead, the oscillations induce fermion particle production and flux propagation, suggesting that the fermion-soliton system does not maintain the integrability and stability expected from the purely bosonic sine-Gordon theory. It may be possible in the future to test these theoretical results, in real systems with topological configurations of boson fields coupled to fermionic particles.
Acknowledgments
B. A. is grateful to the late R. Jackiw for many years of helpful discussions and, in particular, introducing him to the topics of solitary waves and fermion fractionalization. The authors also appreciate the assistance of S. Crittenden with the numerical solution of the Dirac equation.